OpenStudy (anonymous):

42 1/2 + 41 1/3 A) 83 5/6 B) 84 2/3 C) 83 2/3 D) 83 2/5

OpenStudy (anonymous):

@Ms-Brains

MsBrains (ms-brains):

Tell me which you think it is? And I will help you :)

OpenStudy (anonymous):

I think its D. @Ms-Brains

MsBrains (ms-brains):

Nope. Try again

OpenStudy (anonymous):

It doesn't ring a bell :(

OpenStudy (anonymous):

83 5/6

MsBrains (ms-brains):

Yes! :)

OpenStudy (unklerhaukus):

\[42\tfrac12 + 41\tfrac13\\ =42+\tfrac12 + 41+\tfrac13\\ =(42+41)+(\tfrac12 +\tfrac13)\\ =(42+41)+(\tfrac36+\tfrac26)\\ =\]

OpenStudy (mathmale):

UncleR has the best approach here. He suggests you add together the integer parts of these mixed numbers: 42+41. That leaves you with the task of combining 1/2 with 1/3. To do that, you'll need to identify and use the LCD. UncleR has identified that correctly also. The LCD is 6. How did he get that? And how did he get the numerators of the final two fractions?